OpenStudy (anonymous):

7x = 2x – 35 what is x?

OpenStudy (calculusxy):

\[\large 7x = 2x - 35\] \[\large 7x - 2x = 2x - 2x - 35\] \[\large 5x = -35\] \[\large \frac{ 5x }{ 5 } = \frac{ -35 }{ 5 }\] \[\large x= -7 \]

OpenStudy (kfireangel):

\[7*x=2*x-35 => x=-7\]